OpenStudy (anonymous):

hey anyone know how to solve t^4-3t^3-10?

myininaya (myininaya):

i don't know how to solve this

myininaya (myininaya):

because there is nothing to solve

OpenStudy (anonymous):

lol i was thinking the same...

OpenStudy (anonymous):

Yes, there's nothing to solve

OpenStudy (anonymous):

firstly equate it to something miss... then take t^2 = m and substitute o get a quadratic nd solve it out

OpenStudy (anonymous):

t^4-3t^3-10 =0 t^4-3t^3=10 t^3(t-3)=10 so t= cube root (10) or t=13

myininaya (myininaya):

@aditya you know its t^4-3t^3-10 that middle term is not t^2 @dhashni thats a crazy method you got there

myininaya (myininaya):

as an incorrect

OpenStudy (anonymous):

thanks for all your help guys this site is great!

OpenStudy (anonymous):

no answer for this ....because by integral root theorem ...

myininaya (myininaya):

there is no equation lol you can't solve this

OpenStudy (anonymous):

i know but i couldn't believe all the replies i got you guys really tried to help me

OpenStudy (anonymous):

hey myininaya ...its all =o

myininaya (myininaya):

i don't see that

myininaya (myininaya):

i just see an expression

myininaya (myininaya):

no equation

OpenStudy (anonymous):

lets assume..sme typo errors..hhehe

myininaya (myininaya):

lol maybe